3p to the 2nd power+2p+2
Answer by shantalboodram(15)

3p^2+2p+2 = 0
Completing the square
p^2+2p/3+2/3 = 0
p^2 + 2p/3 + (1/3)^2 - (1/3)^2 + 2/3 = 0
(p + 1/3)^2 + 2/3 - 1/9 = 0
(p+1/3)^2 = -5/9
p+1/3 = +- i sqr rt (5/9)
p = -1/3 +- i sqr rt (5/9)